An exercise room has six weight-lifting machines that have no motors and seven treadmills, each equipped with a 2.5-hp (shaft output) motor. The motors operate at an average load factor of 0.7, at which their efficiency is 0.77. During peak evening hours, all 13 pieces of exercising equipment are used continuously, and there are also two people doing light exercises while waiting in line for one piece of the equipment. Assuming the average rate of heat dissipation from people in an exercise room is 640 W determine the rate of heat gain of the exercise room from people and the equipment at peak load conditions The rate of heat gain is W.
Answer #1

1 Hp=745.7 Watts

Rate of heat generated by seven treadmills is

Q1=7*(2.5*745.7)*0.7*(1/0.77)

Q1=11683.4 Watts

Rate of heat generated by 15 people

Q2=15*640=9600 Watts

Total rate of heat gain from exercise room is

Q=11683.4+9600 = 21463.4 Watts